a bioengineer preparing cells for cloning bathes a small piece of rat tissue in a tris buffer HO(CH2)3CNH2, the buffer is made by dissolving 58 g of tris (pkb=5.91) with 5.00 g of HCl to make 2.00 L of solution, calculate pH

please go step by step
Answer #1

The moles of the reagents are calculated:

n Tris = m / MM = 58 g / 121.14 g / mol = 0.479 mol

n HCl = 5 g / 36.5 g / mol = 0.137 mol

The HCl reacts with the Tris and forms Tris-HCl, the remaining moles of Tris and the pKa are calculated:

n Tris remaining = 0.479 - 0.137 = 0.342 mol

pKa = 14 - 5.91 = 8.09

The pH is calculated:

pH = pKa + log n Tris / n TrisHCl = 8.09 + log (0.342 / 0.137) = 8.49

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.